OpenStudy (anonymous):

Find the sum of the first 8 terms of the sequence. Show all work for full credit. 1, -3, -7, -11, ...

OpenStudy (anonymous):

Have you studies Arithmetic sequences or progressions?

OpenStudy (anonymous):

@katherineekc ??

OpenStudy (anonymous):

kinda yeah, im just kinda comfused

OpenStudy (anonymous):

Let's find the common difference between the terms -3-(1)=-4 -7-(-3)=-3+3=-4 So it's an Arithmetic Sequence common difference= -4 \[a_0=1\\\ \text{First term}\] Do you get this?

OpenStudy (anonymous):

okay, yes i get that!

OpenStudy (anonymous):

We need to find the 8 th term nth term is given as \[a_n=a_1+(n-1)d\] n=8, d =-4 (common difference) \[a_8=a_1+(8-1)\times d\] \[a_8=1+7\times (-4)\] Can you find the value of a_8

OpenStudy (anonymous):

@katherineekc

OpenStudy (anonymous):

- 32 ?

OpenStudy (anonymous):

\[a_8=1-28\] Now find the value of 8th term

OpenStudy (anonymous):

-27?

OpenStudy (anonymous):

yes now sum of first 8 terms is \[S=\frac {n}{2}\times (a_1+a_8)\] n=8 \[a_1=1\] \[a_8=-27\] Just simple calculation now, can you do it?

OpenStudy (anonymous):

i believe so yes ..

OpenStudy (anonymous):

Could you find the value? take your time :)

OpenStudy (anonymous):

what would the n value b .. just 1?

OpenStudy (anonymous):

n=8

OpenStudy (anonymous):

oh duh sorry.. h.o

OpenStudy (anonymous):

is it -104?

OpenStudy (anonymous):

correct :)

OpenStudy (anonymous):

ok:) thank you so much!!!

OpenStudy (anonymous):

i hate to be a bother, but do you know how to do this problem? Write the sum using summation notation, assuming the suggested pattern continues. 25 + 36 + 49 + 64 + ... + n2 + ...